# LOST-3DSG: Lightweight Open-Vocabulary 3D Scene Graphs with Semantic Tracking in Dynamic Environments

### v1

@@ 1 Introduction
Recent work on scene representation has increasingly adopted 3D Scene Graphs (3DSGs[armeni2019scenegraph, wald2020learning3d] as a flexible and expressive framework for modeling complex environments. By integrating geometric structure with semantic information, 3DSGs describe scenes as collections of object-centric nodes augmented with attributes and relational edges. This abstraction shifts scene understanding from low-level geometry to a structured, object-level representation that explicitly captures entities and their relationships.
@@
Within this line of research, open-vocabulary 3DSGs [gu2024conceptgraphs, koch2024open3dsg, werby2024hierarchical, maggio2024clio] have gained significant attention. These approaches leverage large pre-trained Foundation Models (FMs), including Vision-Language Models (VLMs) and CLIP [radford2021learning], to encode object nodes with rich semantic representations that are not limited to a predefined taxonomy. As a result, robots can recognize and reason about previously unseen objects and concepts, substantially improving generalization in unstructured environments.
@@
Despite these advantages, such expressiveness often comes at a significant computational and memory cost. Many existing methods rely on storing dense CLIP embeddings at the voxel or point level, producing large-scale semantic maps that are expensive to construct, update, and maintain [gu2024conceptgraphs, yan2025dynamic]. These costs are intensified in dynamic environments, where frequent updates are necessary to reflect changes in object pose and state. The repeated extraction, storage, and management of high-dimensional semantic features ultimately limit scalability and real-time applicability, highlighting the need for more efficient representations that preserve open-vocabulary reasoning while reducing computational overhead.
@@
We propose LOST-3DSG, a lightweight open-vocabulary 3D scene graph designed for dynamic environments. In contrast to existing approaches that store dense CLIP embeddings for objects in the scene, LOST-3DSG relies on low-cost word2vec [mikolov2013efficient] embeddings derived from semantic attributes extracted using a VLM. These compact semantic representations enable robust tracking of dynamic objects over time by reasoning at the attribute level. Rather than depending solely on geometric consistency, the system determines whether an observation corresponds to a previously seen object or a new instance by matching its semantic attributes. For example, if an object previously identified as a "red and brown, wooden and metal hammer" reappears at a different location, it is more likely the same object that has moved rather than a newly observed one. By benchmarking our method against CLIP-based approaches and conducting an extensive ablation study, we demonstrate that our system can accurately track objects in the scene while maintaining a low computational footprint. To further validate the proposed approach, we deploy it in a real-world environment using a TIAGo robot [^1] https://pal-robotics.com/robot/tiago/. We summarize our contributions as follows:
@@ 2 Related Work
3D Scene Graphs In recent years, 3DSGs [armeni2019scenegraph, wald2020learning3d, hughes2022hydra] have emerged as a prominent and widely adopted representation for modeling 3D environments. These representations describe a scene as a graph in which objects are encoded as nodes, while semantic, spatial, and functional relationships between them are captured through edges. This structured formulation provides a compact yet expressive way to model the entities present in an environment and to reason about their interactions. A key strength of 3DSGs lies in their flexibility. By tailoring node attributes and relational edges to the requirements of a given application, 3DSGs can support a broad range of downstream robotic tasks, including navigation [yin2024sg], task planning [rana2023sayplan, liu2025delta, musumeci2025context], manipulation [jiang2024roboexp, rotondi2025fungraph, honerkamp2024language], and human-robot interaction [rotondi2025social3dsg, lim2025eventdrivenstorytellingmultiplelifelike]. Moreover, their graph-based structure makes them easily serializable in formats such as JSON, facilitating their integration into LLM-augmented applications [rana2023sayplan, chang2023dlite, gorlo2024trajectory].
@@
Open-Vocabulary Scene Understanding Recent advances in computer vision, together with the development of Large Language Models (LLMs) and other FMs, have enabled open-vocabulary scene understanding and parsing [Zhao_2017_ICCV]. These approaches allow scenes to be analyzed and objects to be detected without being constrained by a fixed, predefined taxonomy of labels, thereby improving generalization to novel and previously unseen concepts. A common strategy to achieve open-vocabulary understanding in 3D environments is to augment existing scene representations with semantic features extracted from pre-trained models such as CLIP [radford2021learning]. By associating CLIP embeddings with elements of a scene representation, whether voxels, points, primitives, or object nodes, it becomes possible to endow a wide range of mapping paradigms with rich, open-vocabulary semantics, including NeRF [engelmann2024opennerf], Gaussian splatting [guo2024semantic], point clouds [Peng_2023_CVPR], and 3DSGs [koch2024open3dsg]. In doing so, however, open-vocabulary 3DSGs such as ConceptGraphs [gu2024conceptgraphs] and DovSG [yan2025dynamic] require storing high-dimensional semantic features for each voxel in the scene, resulting in a queryable representation at the expense of a substantial memory footprint. In contrast, our method enables an open-vocabulary representation while requiring only a fraction of the computational resources used by CLIP-based approaches.
@@
Tracking in Dynamic Environments Object tracking in dynamic environments has long been a fundamental challenge in robotics. Over the years, a wide range of approaches have been proposed, including filtering techniques such as Kalman filters [welch1995introduction], probabilistic formulations [blackman2004multiple], learning-based methods [ondruska2016deep], and generative models [argenziano2025dynamic]. More recently, a growing body of work has focused on semantic tracking, where objects are tracked not only based on their motion but also according to their semantic properties and behaviors. In this direction,  li2024beyond propose tracking entities in video by jointly reasoning about both their trajectories, answering the question of where an object is, and the underlying semantic events, addressing what is happening. Similarly,  zhang2020spatial leverage deep convolutional features extracted from a pre-trained VGG [simonyan2014very] network to continuously track dynamic objects over time. SemTrack [wang2024semtrack] further advances this line of research by introducing the first large-scale dataset designed to train and evaluate models for semantic tracking in unconstrained environments. Inspired by these approaches, we design a semantic tracking algorithm that maintains object identity by reasoning over semantic attributes such as color and material, while updating object positions within the scene graph as the environment evolves. To the best of our knowledge, this is the first method to perform semantic object tracking directly within a 3DSG representation.
@@ 3.1 Perception Module
The PM builds upon the EMPOWER framework [argenziano2024empower], with several modifications tailored to continuous perception in dynamic environments. In particular, we extend EMPOWER with a streaming processing pipeline that incrementally handles sequential observations as the robot navigates through the environment. A first modification concerns open-set object labeling. In the original EMPOWER pipeline, a multi-role planner identifies task-relevant objects, which are then passed to YOLO-World [cheng2024yolo] for detection. Since our objective is not task planning but scene understanding, we replace this component with a direct VLM query. Given an RGB observation $I_{t}$ at time $t$, the VLM produces a set of object labels $\mathcal{L}_{t}=\{\ell_{1},\ell_{2},\dots,\ell_{N_{t}}\},$ corresponding to the objects present in the scene. In our implementation, we use GPT-5-mini as the VLM. The extracted labels are then passed to an open-vocabulary object detector to spatially ground the symbols in the image. Formally, the detector maps the image and label set to a collection of 2D bounding boxes $\mathcal{B}_{t}=\{b_{i}\mid b_{i}=f_{\text{det}}(I_{t},\ell_{i})\},$ where each bounding box $b_{i}$ localizes the corresponding object $\ell_{i}$. Unlike EMPOWER, which relies on YOLO-World, we adopt OWLv2 [minderer2023scaling] due to its superior open-vocabulary detection performance. Each detected bounding box is then processed by EfficientViT-SAM [zhang2024efficientvit] to obtain a pixel-level segmentation mask $m_{i}=f_{\text{seg}}(I_{t},b_{i}),$ which precisely delineates the spatial extent of the object in the image. In parallel, the cropped image region corresponding to each bounding box is analyzed by the VLM to extract a set of semantic attributes $a_{i}=\{\ell,c,m,d\},$ which are later used by the semantic tracking module. Using camera intrinsics $\mathbf{K}$, extrinsics $\mathbf{T}$, and the depth map $D_{t}$, each segmentation mask is reprojected into 3D space to label the point cloud:

### v2

@@ 1 Introduction
Recent work on scene representation has increasingly adopted 3D Scene Graphs (3DSGs[3, 29] as a flexible and expressive framework for modeling complex environments. By integrating geometric structure with semantic information, 3DSGs describe scenes as collections of object-centric nodes augmented with attributes and relational edges. This abstraction shifts scene understanding from low-level geometry to a structured, object-level representation that explicitly captures entities and their relationships.
@@
Within this line of research, open-vocabulary 3DSGs [10, 15, 32, 19] have gained significant attention. These approaches leverage large pre-trained Foundation Models (FMs), including Vision-Language Models (VLMs) and CLIP [25], to encode object nodes with rich semantic representations that are not limited to a predefined taxonomy. As a result, robots can recognize and reason about previously unseen objects and concepts, substantially improving generalization in unstructured environments.
@@
Despite these advantages, such expressiveness often comes at a significant computational and memory cost. Many existing methods rely on storing dense CLIP embeddings at the voxel or point level, producing large-scale semantic maps that are expensive to construct, update, and maintain [10, 33]. These costs are intensified in dynamic environments, where frequent updates are necessary to reflect changes in object pose and state. The repeated extraction, storage, and management of high-dimensional semantic features ultimately limit scalability and real-time applicability, highlighting the need for more efficient representations that preserve open-vocabulary reasoning while reducing computational overhead.
@@
We propose LOST-3DSG, a lightweight open-vocabulary 3D scene graph designed for dynamic environments. In contrast to existing approaches that store dense CLIP embeddings for objects in the scene, LOST-3DSG relies on low-cost word2vec [20] embeddings derived from semantic attributes extracted using a VLM. These compact semantic representations enable robust tracking of dynamic objects over time by reasoning at the attribute level. Rather than depending solely on geometric consistency, the system determines whether an observation corresponds to a previously seen object or a new instance by matching its semantic attributes. For example, if an object previously identified as a "red and brown, wooden and metal hammer" reappears at a different location, it is more likely the same object that has moved rather than a newly observed one. By benchmarking our method against CLIP-based approaches and conducting an extensive ablation study, we demonstrate that our system can accurately track objects in the scene while maintaining a low computational footprint. To further validate the proposed approach, we deploy it in a real-world environment using a TIAGo robot [^1] https://pal-robotics.com/robot/tiago/. We summarize our contributions as follows:
@@ 2 Related Work
3D Scene Graphs In recent years, 3DSGs [3, 29, 13] have emerged as a prominent and widely adopted representation for modeling 3D environments. These representations describe a scene as a graph in which objects are encoded as nodes, while semantic, spatial, and functional relationships between them are captured through edges. This structured formulation provides a compact yet expressive way to model the entities present in an environment and to reason about their interactions. A key strength of 3DSGs lies in their flexibility. By tailoring node attributes and relational edges to the requirements of a given application, 3DSGs can support a broad range of downstream robotic tasks, including navigation [34], task planning [26, 18, 22], manipulation [14, 27, 12], and human-robot interaction [4, 17]. Moreover, their graph-based structure makes them easily serializable in formats such as JSON, facilitating their integration into LLM-augmented applications [26, 6, 9].
@@
Open-Vocabulary Scene Understanding Recent advances in computer vision, together with the development of Large Language Models (LLMs) and other FMs, have enabled open-vocabulary scene understanding and parsing [37]. These approaches allow scenes to be analyzed and objects to be detected without being constrained by a fixed, predefined taxonomy of labels, thereby improving generalization to novel and previously unseen concepts. A common strategy to achieve open-vocabulary understanding in 3D environments is to augment existing scene representations with semantic features extracted from pre-trained models such as CLIP [25]. By associating CLIP embeddings with elements of a scene representation, whether voxels, points, primitives, or object nodes, it becomes possible to endow a wide range of mapping paradigms with rich, open-vocabulary semantics, including NeRF [8], Gaussian splatting [11], point clouds [24], and 3DSGs [15]. In doing so, however, open-vocabulary 3DSGs such as ConceptGraphs [10] and DovSG [33] require storing high-dimensional semantic features for each voxel in the scene, resulting in a queryable representation at the expense of a substantial memory footprint. In contrast, our method enables an open-vocabulary representation while requiring only a fraction of the computational resources used by CLIP-based approaches.
@@
Tracking in Dynamic Environments Object tracking in dynamic environments has long been a fundamental challenge in robotics. Over the years, a wide range of approaches have been proposed, including filtering techniques such as Kalman filters [31], probabilistic formulations [5], learning-based methods [23], and generative models [2]. More recently, a growing body of work has focused on semantic tracking, where objects are tracked not only based on their motion but also according to their semantic properties and behaviors. In this direction,  Li et al. [16] propose tracking entities in video by jointly reasoning about both their trajectories, answering the question of where an object is, and the underlying semantic events, addressing what is happening. Similarly,  Zhang et al. [35] leverage deep convolutional features extracted from a pre-trained VGG [28] network to continuously track dynamic objects over time. SemTrack [30] further advances this line of research by introducing the first large-scale dataset designed to train and evaluate models for semantic tracking in unconstrained environments. Inspired by these approaches, we design a semantic tracking algorithm that maintains object identity by reasoning over semantic attributes such as color and material, while updating object positions within the scene graph as the environment evolves. To the best of our knowledge, this is the first method to perform semantic object tracking directly within a 3DSG representation.
@@ 3.1 Perception Module
The PM builds upon the EMPOWER framework [1], with several modifications tailored to continuous perception in dynamic environments. In particular, we extend EMPOWER with a streaming processing pipeline that incrementally handles sequential observations as the robot navigates through the environment. A first modification concerns open-set object labeling. In the original EMPOWER pipeline, a multi-role planner identifies task-relevant objects, which are then passed to YOLO-World [7] for detection. Since our objective is not task planning but scene understanding, we replace this component with a direct VLM query. Given an RGB observation $I_{t}$ at time $t$, the VLM produces a set of object labels $\mathcal{L}_{t}=\{\ell_{1},\ell_{2},\dots,\ell_{N_{t}}\},$ corresponding to the objects present in the scene. In our implementation, we use GPT-5-mini as the VLM. The extracted labels are then passed to an open-vocabulary object detector to spatially ground the symbols in the image. Formally, the detector maps the image and label set to a collection of 2D bounding boxes $\mathcal{B}_{t}=\{b_{i}\mid b_{i}=f_{\text{det}}(I_{t},\ell_{i})\},$ where each bounding box $b_{i}$ localizes the corresponding object $\ell_{i}$. Unlike EMPOWER, which relies on YOLO-World, we adopt OWLv2 [21] due to its superior open-vocabulary detection performance. Each detected bounding box is then processed by EfficientViT-SAM [36] to obtain a pixel-level segmentation mask $m_{i}=f_{\text{seg}}(I_{t},b_{i}),$ which precisely delineates the spatial extent of the object in the image. In parallel, the cropped image region corresponding to each bounding box is analyzed by the VLM to extract a set of semantic attributes $a_{i}=\{\ell,c,m,d\},$ which are later used by the semantic tracking module. Using camera intrinsics $\mathbf{K}$, extrinsics $\mathbf{T}$, and the depth map $D_{t}$, each segmentation mask is reprojected into 3D space to label the point cloud: